OpenStudy (anonymous):

Solve for x \[\huge (x ^{2}-9)(\left| x \right|-2) \le 0 \]

OpenStudy (triciaal):

do you recognize the difference of 2 squares? x^2 - 9 = (x + 3)(x-3) [x] means use the positive value use this information in the original (x-3)(x+3)(x-2) < = 0

OpenStudy (triciaal):

values of x when =0 are 3, -3, 2

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then plot on number line ?